lkyl Fliers Company produces model airplanes. During the month of November, it produced 3,000 planes. The actual labor hours were 5 hours per plane. Its standard labor hours are 10 hours per plane. The standard labor rate is $8 per hour. At the end of November, Alkyl Fliers found that it had a favorable labor rate variance of $12,000. What was Alkyl Fliers' actual cost per labor hour? a. $7.60 b. $9.50 c. $12.75 d. $7.20
Answer #1

Labor rate variance = (SR-AR) * AH

(8 - AR) * (3000*5) = 12000

(8-AR) = 12000/15000 = 0.8

AR = 8 - 0.8 = 7.20

OPTION D IS THE ANSWER